I must decide if the algebraic manipulation is correct or incorrect. if it is incorrect I need to fix it or explaine why it cannot be simplified.I don't know if this is the right spot to post but if you can help that would be great.
3^x+3^x+3^x=3^3^x

Answer by ankor@dixie-net.com(12693)

3^x+3^x+3^x = 3^3^x
:
not right, think of it like this:
1(3^x) + 1(3^x)+ 1(3^x) = 3(3^x)